# Interpreting the Hi cm cosmology maps through Largest Cluster Statistics. Part II. Impact of the realistic foreground and instrumental noise on synthetic SKA1-Low observations

### v1

@@ Abstract
The Largest Cluster Statistics (LCS) analysis of the redshifted 21 cm maps has been demonstrated to be an efficient and robust method for following the time evolution of the largest ionized regions (LIRs) during the Epoch of Reionization (EoR). The LCS can, in principle, constrain the reionization model and history by quantifying the morphology of neutral hydrogen (Hi) distribution during the different stages of the EoR. Specifically, the percolation transition of ionized regions, quantified and constrained via LCS, provides a crucial insight about the underlying reionization model. The previous LCS analysis of EoR 21 cm maps demonstrates that the convolution of the synthesized beam of the radio interferometric arrays, e.g. SKA1-Low with the target signal, shifts the apparent percolation transition of ionized regions towards the lower redshifts. In this study, we present an optimal thresholding strategy to reduce this bias in the recovered percolation transition. We assess the robustness of LCS analysis of the 21 cm maps in the presence of antenna-based gain calibration errors and instrumental noise for SKA1-Low. This analysis is performed using synthetic observations simulated by the 21cmE2E pipeline, considering SKA1-Low AA4 configuration within a radius of 2 km from the array centre. Our findings suggest that a minimum of hours of observation (SNR ) are required for the LCS analysis to credibly suppress the confusion introduced by thermal noise. Further, we also demonstrate that for a maximum antenna-based calibration error tolerance of (post calibration), the reionization history can be recovered in a robust and relatively unbiased manner using the LCS.
@@ 1 Introduction
Motivated by this, we present an optimal thresholding strategy to reduce this bias in the recovered percolation transition in this study. Furthermore, we assess the robustness of our LCS analysis of the 21 cm maps under various kinds of foreground corruption conditions. For this purpose, we consider two sources of foreground contamination: a) extra-galactic point sources and b) diffuse synchrotron and free-free emission. For all of these sky models, we consider the antenna-based gain calibration errors to estimate the maximum tolerance level to recover the reionization history in a robust and relatively unbiased manner using LCS. In addition, we have also investigated the impact of instrumental noise for SKA1-Low on this analysis via the synthetic observations simulated by the 21cmE2E pipeline 222https://gitlab.com/samit-pal/21cme2e. This analysis is done by diagnostic tool SURFGEN2 [29, 30] to estimate the LCS and gain insights into the percolation of Hii regions during the EoR.
@@ 2.2 Astrophysical foregrounds
One of the major contaminants of CD/EoR experiments is the astrophysical foreground. Their brightness is orders of magnitude higher than the faint cm signal. The primary components in the foreground include diffuse galactic synchrotron radiation, galactic and extra-galactic free-free radiation and extra-galactic point sources. The foreground emission is expected to be spectrally smooth. However, calibration and instrumental effects introduce additional unsmooth structures. We considered two sources of foreground contamination: diffuse emission and extra-galactic point sources. The foreground contributions are detailed below.
@@ 2.2.2 Extragalactic point sources
In addition to the galactic and extra-galactic diffuse emission, the total foreground emission is also affected by extra-galactic point sources. These extra-galactic point sources are often compact and finite in size. This study uses the Tiered Radio Extra-galactic Continuum Simulation (T-RECS)[47] catalogue. At a frequency of MHz, the flux values of the sources range from mJy to Jy. The fluxes were transformed to extrapolate to the frequencies of interest using the relationship , where is . These extra-galactic point sources mainly comprises of star-forming galaxies and radio-quiet quasars. For a detailed description of the T-RECS catalogue model, readers can refer to [47, 16]. The T-RECS catalogue covers the sky area. However, we generated image cubes only for the central region, covering fields. This was done to match the field of view (FoV) of the input of the input Hi maps from reionyuga, depending on the redshift of observation. Bright sources in the beam sidelobes present challenges for data calibration in real observations. However, these effects were not considered in this work.
@@ 3 Simulations of observations
This section outlines the synthetic observation with SKA1-Low to simulate the radio sky discussed in Section 2. Figure 1 shows the schematic diagram of the 21cmE2E pipeline. This pipeline is based on the OSKAR555https://ska-telescope.gitlab.io/sim/oskar/ and CASA666https://casa.nrao.edu/ [48]. The OSKAR [49] package generates the simulated visibility based on the input sky model, observational parameters, and telescope specifications. CASA is then used to create image cubes from the simulated visibilities. We simulated only minutes of observation, which will be restricted to minutes around the transit time of the target EoR field for this study to reduce computational costs. The field of view centred at h and . The integration time for the simulation was set to 120 seconds. In the simulated Hi cm cube, one axis represents the line-of-sight (LoS) or frequency axis. The cubes were then divided into slices based on their frequency resolution. Each slice was assigned a frequency label corresponding to the comoving distance from the observer. The slices of the cm maps were converted from comoving Mpc (cMpc) to angular coordinates on the sky plane. These angular maps were then processed through the 21cmE2E pipeline. Images were produced using the briggs weighting scheme with robust parameter using CASA tclean task. The frequency-labeled slices were stacked, and final image cubes were made for further analysis. It is noted that the observed Hi maps have the same size as the input Hi maps. The pixel size was chosen to match the size of the input Hi maps.
@@ 3.2 Gain calibration error
Where & is the complex antenna gain term and is the noise on th baseline.However, the true visibility, , can be measured by a perfectly calibrated noiseless interferometer. To obtain the true visibility, we need to calibrate the observation. During calibration, we solve the gain factors & to minimize the chi-square between true and observed visibility. However, the accuracy of calibration is limited by the SNR. In an ideal scenario, the gain factor should be unity. However, an analytical solution of Equation 3.1 for the desired parameter is intractable. Given antennas, there are baselines. These baselines correspond to true visibilities, which we want to solve along with antenna gain factors. This process of correcting the antenna gain is known as self-calibration or direction-independent calibration. In cm experiments, either sky-based or redundant calibration techniques are used. The accurate calibration is essential because the uncalibrated part of gains or residual gains will propagate into the subsequent steps, thereby introducing errors during the extraction of the target signal. The residual complex gain from the antenna ‘i’ can be modelled by
@@
In order to add the calibration error in this simulation, we assumed that the errors in amplitude and phase follow a zero-mean Gaussian distribution with a specified standard deviation (see equation 3.2). The calibration errors are sampled randomly with different standard deviations at the level of to in amplitude and to degree in phase. These errors are applied to each time domain of a single observation. We assumed that the calibration errors were uncorrelated for multiple observations. Therefore, calibration errors can be suppressed by averaging the deep observations. We focus on the deep observations with a total duration of h. In this work, we simulated the input sky with min of continuous track observation. Therefore, to achieve the h of integration time, the same patch of sky must be observed repeatedly times. After the final post-calibration and averaging, the standard deviations are calculated for each time domain of a single observation. The calibration error is multiplied by the model visibility data. This creates the residual visibility after subtracting the foreground models using CASA uvsub task. This uvsubed residual visibilities dataset is used to make image cubes for LCS analysis. The top panel of Figure 3 shows the slices of the observed brightness temperature of ). The top-left panel shows the observed Hi map without any bias due to corruption. The middle and right top panel shows the residual maps after introducing calibration errors of & , respectively. The bottom panel shows the recovered Hii regions identified using the triangle algorithm. In these maps, the white (black) regions represent the ionized (neutral) regions. The higher calibration errors reduce the dynamic range of the image. Residual foregrounds pose a significant challenge in making the map of ionized regions. At the calibration inaccuracy of , the triangle method failed to retrieve the regions from the residual field.
@@ 3.3 Instrumental noise
Where is the system temperature, is the effective area of the antenna/station, is the Boltzmann constant, is the frequency resolution and is the integration time of the visibilities. The 777https://www.skao.int/sites/default/files/documents/SKAO-TEL-0000818-V2_SKA1_Science_Performance.pdf values for SKA1-Low are listed by Braun et al. [53]. We interpolated the values to the frequencies of interest. For this case, we assume the residual foreground contaminations are below the Hi signal level. The only contamination present in the Hi maps is thermal noise. The contribution of the thermal noise in the visibility domain is rescaled by the factor of , where h represents the observation time per day will be restricted to minutes around the transit time of the target EoR field and is the total integration time. This rescaling represents the coherent averaging of visibility data over the total integration time. For this simulation, we vary the total integration time of , , and 888Our simulations track the sky for only 30 minutes per observation for computational efficiency. To accumulate hours of observation time, this requires repetitions. In contrast, actual observations with a 4 hour daily tracking time would necessitate repetitions of the same sky patch, taking approximately two years to achieve the desired observation time. If we shift the phase centre a bit of arcsec, we can mitigate the systematics effect. to achieve a good signal-to-noise ratio level (SNR ) in the synthesized map.
@@
To incorporate the thermal noise into the simulation, we downsampled the simulated 21 cm maps by a factor of along each side. As discussed in our previous study, the synthetic cm observations require input maps with dimensions of grids. We used a downsampling algorithm available in the python package scikit-image and integrated this method into the 21cmE2E pipeline. This downsampling is essential to reduce the total observation time to achieve the desired SNR level. After downsampling the maps, the grid size of the final maps is cMpc for redshift , which corresponds to an angular resolution of arcmin and a frequency resolution of MHz. We generated the image cubes using the 21cmE2E pipeline after adding the thermal noise. In order to balance the resolution and the sensitivity of the telescope, two different imaging weighting schemes are used: natural weighting and briggs weighting. The natural weighting scheme provides a relatively higher SNR but a lower resolution compared to the briggs weighting scheme with a robust parameter of . The left and middle panel of Figure 4 shows the slices of the brightness temperature field before and after adding the thermal noise with an observation time of h at a channel width of MHz, respectively, at neutral fraction (). It is seen that more tunnel-like artefacts are visible on the map after introducing thermal noise. These artefacts result from the effect of noise and deconvolution. In the later section 5.3, we discuss how this feature will affect the percolation process. The right panel of Figure 4 shows the recovered Hii regions after applying an entropy-based thresholding algorithm (Yen’s method). In this map, white regions represent recovered Hii regions, while black regions represent Hi regions.
@@ 4.1 Binarization of the image cubes
In order to estimate the LCS on the brightness temperature maps, we use a code, SURFGEN2 [60, 61, 29, 30]. SURFGEN2 not only determines the LCS but also helps identify the topological and morphological features of each ionized region using Shapefinders [62]. 999Shapefinders are derived from the ratios of the Minkowski functionals. For instance, in three dimensions, the four Minkowski functionals give rise to three shapefinders, with each one representing the extent of a closed surface along one of the three dimensions; see [62] for more details. For an in-depth understanding of the operating principles of the SURFGEN2 code, readers can refer to [61, 30]. The SURFGEN2 code requires thresholding to binarize the neutral and ionized regions of the Hi maps. In an ideal Hi 21 cm brightness temperature field (), an ionized region is identified by equal to zero. For an interferometric observation, we measure only the fluctuations in the cm signal. Accordingly, the minima of the brightness temperature maps correspond to ionized regions, as expected. However, in real observations, the presence of systematic effects poses a challenge in finding the optimal threshold to distinguish the ionized regions. In our previous work, Dasgupta et al. [33], imposed a gradient-descent method on the histograms of the synthetically observed Hi maps by the SKA1-Low to binarize them. When we add calibration errors and instrumental noise to the Hi maps, finding an optimal threshold becomes a challenging task. In this paper, we employ the intensity-based thresholding algorithm (i.e., triangle method [63, 64]) for calibration errors and the entropy-based thresholding algorithm (i.e., Yen’s method) [65, 64] for instrumental noise to separate image pixels. In future work, we will investigate the optimal thresholding for binarizing the Hi field from interferometric observations (Dasgupta et. al., in prep).

### v2

@@ Abstract
The Largest Cluster Statistics (LCS) analysis of the redshifted 21 cm maps has been demonstrated to be an efficient and robust method for following the time evolution of the largest ionized regions (LIRs) during the Epoch of Reionization (EoR). The LCS can, in principle, constrain the reionization model and history by quantifying the morphology of neutral hydrogen (Hi) distribution during the different stages of the EoR. Specifically, the percolation transition of ionized regions, quantified and constrained via LCS, provides a crucial insight about the underlying reionization model. The previous LCS analysis of EoR 21 cm maps demonstrates that the convolution of the synthesized beam of the radio interferometric arrays, e.g. SKA1-Low with the target signal, shifts the apparent percolation transition of ionized regions towards the lower redshifts. In this study, we present an optimal thresholding strategy to reduce this bias in the recovered percolation transition. We assess the robustness of LCS analysis of the 21 cm maps, considering the effects of antenna-based gain calibration errors and instrumental noise for SKA1-Low. This analysis is performed using synthetic observations simulated by the 21cmE2E pipeline, considering SKA1-Low AA4 configuration within a radius of 2 km from the array centre. Our findings suggest that a minimum of hours of observation (SNR ) are required for the LCS analysis to credibly suppress the confusion introduced by thermal noise. Further, we also demonstrate that for a maximum antenna-based calibration error tolerance of (post calibration), the reionization history can be recovered in a robust and relatively unbiased manner using the LCS.
@@ 1 Introduction
Motivated by this, we present an optimal thresholding strategy to minimize the bias in the recovered percolation transition. We further assess the robustness of our LCS analysis of the 21 cm maps under various foreground contamination scenarios. For this purpose, we consider two sources of foreground contamination: a) extra-galactic point sources and b) diffuse synchrotron and free-free emission. For all of these sky models, we consider the antenna-based gain calibration errors to estimate the maximum tolerance level to recover the reionization history in a robust and relatively unbiased manner using LCS. In addition, we have also investigated the impact of instrumental noise for SKA1-Low on this analysis via the synthetic observations simulated by the 21cmE2E pipeline222https://gitlab.com/samit-pal/21cme2e. This analysis is done by diagnostic tool SURFGEN2 [29, 30] to estimate the LCS and gain insights into the percolation of Hii regions during the EoR.
@@ 2.2 Astrophysical foregrounds
One of the major contaminants of CD/EoR experiments is the astrophysical foreground. Their brightness is orders of magnitude higher than the faint cm signal. The primary components in the foreground include diffuse galactic synchrotron radiation, galactic and extra-galactic free-free radiation and extra-galactic point sources. The foreground emission is expected to be spectrally smooth. However, calibration and instrumental effects introduce additional unsmooth structures. We considered two sources of foreground contamination, diffuse emission and extra-galactic point sources, to test these impacts on the residual maps. The foreground contributions are detailed below.
@@ 2.2.2 Extragalactic point sources
In addition to the galactic and extra-galactic diffuse emission, the total foreground emission is also affected by extra-galactic point sources. These extra-galactic point sources are often compact and finite in size. This study uses the Tiered Radio Extra-galactic Continuum Simulation (T-RECS)[47] catalogue. At a frequency of MHz, the flux values of the sources range from mJy to Jy. The fluxes were transformed to extrapolate to the frequencies of interest using the relationship , where is . These extra-galactic point sources mainly comprises of star-forming galaxies and radio-quiet quasars. For a detailed description of the T-RECS catalogue model, readers can refer to [47, 16]. The T-RECS catalogue comprises of 2522 unpolarized flat-spectrum sources within the sky area. However, we generated image cubes only for the central region, covering fields. This was done to match the field of view (FoV) of the input of the input Hi maps from reionyuga, depending on the redshift of observation. Bright sources in the beam sidelobes present challenges for data calibration in real observations. However, these effects were not considered in this work.
@@ 3 Simulations of observations
This section outlines the synthetic observation with SKA1-Low to simulate the radio sky discussed in Section 2. The observation parameters are listed in table 1. Figure 1 shows the schematic diagram of the 21cmE2E pipeline. This pipeline is based on the OSKAR555https://ska-telescope.gitlab.io/sim/oskar/. The OSKAR [48] package generates the simulated visibility based on the input sky model, observational parameters, and telescope specifications. We simulated only hours of observation, which will be restricted to hours around the transit time of the target EoR field for this study to reduce computational costs. The field of view centred at h and . The integration time for the simulation was set to 10 seconds. In the simulated Hi cm cube, one axis represents the line-of-sight (LoS) or frequency axis. The cubes were then divided into slices based on their frequency resolution. Each slice was assigned a frequency label corresponding to the comoving distance from the observer. The slices of the cm maps were converted from comoving Mpc (cMpc) to angular coordinates on the sky plane. These angular maps were then processed through the 21cmE2E pipeline. We use wsclean to produce the image cube from simulated visibilities using the Briggs weighting scheme with robust parameter . The frequency-labeled slices were stacked, and final image cubes were made for further analysis. It is noted that the observed Hi maps have the same size as the input Hi maps. The pixel size was chosen to match the size of the input Hi maps.
| Parameter | Value | |
| Redshift range | 7.2 - 8.8 | |
| Telescope | SKA1-Low AA4 | |
| Number of stations | 296 | |
| Maximum baseline | km | |
| Polarization | Stokes I | |
| Phase Center(J2000) | RA, DEC= 15 h, | |
| Duration of obseravtion | 2 h ( HA) | |
| Time resolution | 10 s | |
| Sky model | 2522 point sources in the central |
@@ 3.2 Gain calibration error
Where & is the complex antenna gain term and is the noise on th baseline. However, the true visibility, , can be measured by a perfectly calibrated noiseless interferometer. To obtain the true visibility, we need to calibrate the observation. During calibration, we solve the gain factors & to minimize the chi-square between true and observed visibility. However, the accuracy of calibration is limited by the SNR. In an ideal scenario, the gain factor should be unity. However, an analytical solution of Equation 3.1 for the desired parameter is intractable. Given antennas, there are baselines. These baselines correspond to true visibilities, which we want to solve along with antenna gain factors. This process of correcting the antenna gain is known as self-calibration or direction-independent calibration. In cm experiments, either sky-based or redundant calibration techniques are used. The accurate calibration is essential because the uncalibrated part of gains or time and frequency-correlated residual gains will propagate into the subsequent steps, thereby confusing or even obscuring during the extraction of the target signal. The complex gain from the antenna ‘i’ can be modelled by
@@
To simulate antenna-based gain calibration errors, we added a noise term to the gain model with a specified standard deviation in amplitude and phase for both time and frequency domains (see equation 3.2). The calibration errors are sampled randomly with different standard deviations at the level in the range from to in amplitude and from to degree in phase. The amplitude and phase response of the residual gain errors value for each station changes at every second time stamp and every frequency channel throughout the observation. The mean amplitude response is set to unity and the mean phase response to zero, ensuring no systematic calibration offsets are introduced. To simulate realistic calibration errors, we assume that time and frequency fluctuations are independent, and the gain error at each time and frequency point is given by the complex product of two one-dimensional distributions- one for time and one for frequency. These one-dimensional gain distributions are initially populated with random samples from a Gaussian distribution, with specified standard deviations for both amplitude and phase. We then apply the colorednoise666https://github.com/felixpatzelt/colorednoise.git Python code to transform these white-noise distributions, which have equal power at all scales, into red-noise distributions with a power-law index. This process enhances the fluctuation power on the longest timescales and across the broadest frequency ranges while preserving the original mean and standard deviation. These final gain errors closely resemble the correlated error patterns observed in real calibration data. This approach to simulating direction-independent calibration errors was introduced in the SKA Science Data Challenge 3a (SDC3a) [51]. The gain correlations are usually at time scales of a few tens of seconds [52, 53]. Since any such calibration procedure is done over short time scales and narrow frequency bands, we assume that residual calibration errors are not correlated beyond each night/day’s observations (i.e. 2 hours). So, any systematic errors are only restricted within the 2 hours of observing time. Hence, the RMS noise floor achieved after each epoch (2 hours) of observations is then co-added with other epochs, and the RMS noise reduces as , where is the number of independent observing epochs. The assumptions made in our study are consistent with the previous work by Datta et al. [54]. We focus on deep observations with a total duration of 1000 hours. Our simulation models this deep integration by assuming 500 times repetition of multiple nights of continuous 2-hour tracking observations of the same patch. The final post-calibration and post-averaging standard deviation values are then applied to each corresponding single observation period. This work does not apply any mitigation techniques, as the goal is to estimate the accuracy of any such techniques that will allow us to detect the EoR signal.
The calibration errors were applied by multiplying them with the model visibility data. The residual visibilities were then obtained by subtracting the foreground models using CASA uvsub task. This uvsubed residual visibilities dataset is used to make image cubes for LCS analysis. The top panel of Figure 3 shows the slices of the observed brightness temperature map at ). The top-left panel shows the observed Hi map without any bias due to corruption. The middle and right top panel shows the residual maps after introducing calibration errors of & , respectively. The bottom panel shows the recovered Hii regions identified using the optimum thresholding method. The white (black) regions in these maps represent the recovered neutral (ionized) regions. At the higher calibration errors level (calibration inaccuracy of ), residual foreground contamination introduces artificial filamentary or tunnel-like features (deconvolution artefacts) into the obtained 21 cm observation images.
@@ 3.3 Instrumental noise
Where is the system temperature, is the effective area of the antenna/station, is the Boltzmann constant, is the frequency resolution and is the integration time of the visibilities. The 777https://www.skao.int/sites/default/files/documents/SKAO-TEL-0000818-V2_SKA1_Science_Performance.pdf values for SKA1-Low are listed by Braun et al. [56]. We interpolated the values to the frequencies of interest. Our previous study by Dasgupta et al. [33] introduced additive Gaussian noise in the image plane. The value of systematic noise is drawn from a zero-mean Gaussian random distribution with a specified standard deviation and added to the pixel values of an image. In contrast, this work uses a more realistic representation of instrumental noise in radio interferometric observations by introducing an uncorrelated Gaussian noise in the visibility domain. Furthermore, while Dasgupta et al. [33] applied a constant RMS noise value to the image, our method accounts for the system noise RMS as a function of observing frequency, reflecting the actual variation of noise characteristics across the observational bandwidth. For this case, we assume the residual foreground contaminations are below the Hi signal level. The only contamination present in the Hi maps is thermal noise. The contribution of the thermal noise in the visibility domain is rescaled by the factor of , where h represents the observation time per day will be restricted to hours around the transit time of the target EoR field and is the total integration time. This rescaling represents the coherent averaging of visibility data over the total integration time. For this simulation, we vary the total integration time of , , and 888Our simulations track the sky for only 2 hours per observation for computational efficiency. To accumulate hours of observation time, this requires repetitions. In contrast, actual observations with a 4 hour daily tracking time would necessitate repetitions of the same sky patch, taking approximately two years to achieve the desired observation time. If we shift the phase centre a bit of arcsec, we can mitigate the systematics effect. to achieve a good signal-to-noise ratio level (SNR ) in the synthesized map. The SNR of the image is defined as the ratio of the rms fluctuation of the target signal in the image cube to the rms thermal noise limit for SKA1-Low, evaluated over the total observation time and on the scale of the angular resolution element ().
@@
To incorporate the thermal noise into the simulation, we downsampled the simulated 21 cm maps by a factor of along each side. As discussed in our previous study, the synthetic cm observations require input maps with dimensions of grids. We used a downsampling algorithm available in the python package scikit-image and integrated this method into the 21cmE2E pipeline. This downsampling is essential to reduce the total observation time to achieve the desired SNR level. After downsampling the maps, the grid size of the final maps is cMpc for redshift , which corresponds to an angular resolution of arcmin and a frequency resolution of MHz. We generated the image cubes using the 21cmE2E pipeline after adding the thermal noise. In order to balance the resolution and the sensitivity of the telescope, two different imaging weighting schemes are used: natural weighting and briggs weighting. The natural weighting scheme provides a relatively higher SNR but a lower resolution compared to the briggs weighting scheme with a robust parameter of . The left and middle panel of Figure 4 shows the slices of the brightness temperature field before and after adding the thermal noise with an observation time of h at a channel width of MHz, respectively, at neutral fraction . It is seen that more tunnel-like artefacts are visible on the map after introducing thermal noise. These artefacts result from the effect of noise and deconvolution. In the later section 5.3, we discuss how this feature will affect the percolation process. The right panel of Figure 4 shows the recovered Hii regions after applying an optimum thresholding method. In this map, black(white) regions represent recovered ionized (neutral) regions.
@@ 4.1 Binarization of the image cubes
In order to estimate the LCS on the brightness temperature maps, we use a code, SURFGEN2 [63, 64, 29, 30]. SURFGEN2 not only determines the LCS but also helps identify the topological and morphological features of each ionized region using Shapefinders [65]. 999Shapefinders are derived from the ratios of the Minkowski functionals. For instance, in three dimensions, the four Minkowski functionals give rise to three shapefinders, with each one representing the extent of a closed surface along one of the three dimensions; see [65] for more details. For an in-depth understanding of the operating principles of the SURFGEN2 code, readers can refer to [64, 30]. The SURFGEN2 code requires thresholding to binarize the neutral and ionized regions of the Hi maps. In an ideal Hi 21 cm brightness temperature field (), an ionized region is identified by equal to zero. For an interferometric observation, we measure only the fluctuations in the cm signal. Accordingly, the minima of the brightness temperature maps correspond to ionized regions, as expected. However, in real observations, the presence of systematic effects poses a challenge in finding the optimal threshold to distinguish the ionized regions. In our previous work, Dasgupta et al. [33], imposed a gradient-descent method on the histograms of the synthetically observed Hi maps by the SKA1-Low to binarize them. When we add calibration errors and instrumental noise to the Hi maps, finding an optimal threshold becomes a challenging task.
In this work, we developed a modified approach to identify ionized regions in the 21 cm observation maps by combining a global thresholding method with the unsharp masking technique. Unsharp masking is a widely used image enhancement method in image processing that sharpens features by emphasizing edges and fine details. It is a linear operation, making it a preferred choice over deconvolution, which is often an ill-posed problem. This technique enhances edges (and other small-scale features in an image) by subtracting an unsharp or smoothed (blurred) version of the image from the original, thereby highlighting small-scale structures. The unsharp masking operation is mathematically expressed as:
where is the original image, is its smoothed version, and is a scaling factor that controls the strength of the enhancement. A higher value of increases the contrast of fine details, while a lower value results in more subtle sharpening. We applied the unsharp masking algorithm using the implementation from the scikit-image Python package. This step better defines the boundaries of ionized bubbles prior to applying a global thresholding criterion for more precise segmentation. We apply the method of optimum thresholding to binarize the observed 21 cm maps produced by the 21cmE2E-pipeline and calculate LCS for each of the image cubes.
To assess the effectiveness of our method, we compare it with the binarization technique described in Giri et al. [66]. We observed that their approach tends to over-segment ionized regions in the presence of low-density fluctuations within neutral regions. However, our method demonstrates robustness across diverse density environments and more accurately preserves the morphological structure of ionized regions (see Appendix A). It is important to note that the binarization technique can be biased by the instrumental noise. The instrumental noise introduces the small-scale features in the HI maps from 21 cm observation. In order to mitigate this, a Gaussian filter is applied prior to our modified binarization scheme. This step ensures robust segmentation from noisy 21 cm observation data. This method, herein termed optimal thresholding, is proposed as the robust thresholding method in all scenarios. In future work, we will also investigate the optimal thresholding for binarizing the Hi field from interferometric observations (Dasgupta et. al., in prep).